OpenStudy (anonymous):

How do you factor 25x^3 - 15x^2 |||| Would the answer be just 5x-3?

OpenStudy (skullpatrol):

To factor $$\Huge25x^3 - 15x^2 $$ you first look for the greatest common factor (GCF) of the coefficients which is the greatest integer that is a factor of both of them i.e. the product of the smaller powers of each common prime factor.$$25 = 5*5$$ $$15 = 3*5$$which in this case is 5. Next you find the greatest common factor of the two monomials which is the common factor with the greatest coefficient, 5 in this case, and greatest degree in each variable, in this case. $$x^2$$ This gives the factored form: $$\Huge 5x^2(5x-3)$$ Be sure to check your work by using the distributive property: $$\Huge \underbrace{a(b-c)}=ab-ac$$ factored form